# Amazon Web Service

### v1

@@ 2.4 Compensatory Strategies and Architectural Remedies
Tool-augmented approaches explicitly address the execution gap. Toolformer (Schick et al., 2023) enables automated tool calling, while ReAct (Yao et al., 2022) combines reasoning with external action. Yang et al. (Yang et al., 2024) provide empirical validation that this separation—LLMs for predicate extraction, external systems for computation—consistently outperforms end-to-end LLM execution across arithmetic tasks.
@@ 3 Structural Limits on Symbolic Computation
The computational split-brain syndrome manifests most clearly in symbolic computation, where LLMs can perfectly explain algorithms they cannot reliably execute. This systematic dissociation—drawing analogy to neurological conditions where different brain systems cannot coordinate effectively—emerges as geometric separation in representational space: models develop distinct pathways for “knowing about" procedures versus “executing” them.
@@ 3.1 Claim 1: Contextual Averaging Prevents Automatic Domain Binding
where $C(w)$ denotes all contexts containing token $w$. This produces context-averaged embeddings that trade off among multiple usage modes, similar to distributional semantics models like Word2Vec Mikolov et al. (2013).
@@
These geometric failures directly impede the formation of stable arithmetic circuits that could operate reliably across different input formats or numerical ranges. McLeish et al. (McLeish et al., 2024) demonstrate that altering input geometry alone can dramatically improve simple numerical reasoning, which might appear to contradict our contextual averaging argument. However, their approach involves training models exclusively on arithmetic problems, which inadvertently avoids the fundamental contextual averaging problem faced by real-world LLMs. Once embeddings are contaminated through multi-domain training—where tokens like “9.11” accumulate conflicting associations across historical, versioning, and mathematical contexts—this geometric solution breaks down. The McLeish results thus support rather than refute our analysis: they show embedding geometry matters precisely because real-world training makes it unsolvable.
@@ 3.2 Claim 2: Feed-Forward Networks Resort to Pattern Storage
Given contextually-averaged symbolic representations (Claim 1), feed-forward networks face an additional fundamental constraint: the architectural impossibility of implementing exact symbolic operations through weight configuration alone. With direct computation blocked, FFNs resort to residual fitting by memorizing patterns rather than implementing symbolic procedures, despite their theoretical computational capacity.
The Computational Burden Falls on FFNs.
The transformer architecture creates a clear computational division. While attention mechanisms perform complex contextual computations—processing relationships between operands through learned weighted combinations and integrating information across the evolving sequence representation—they operate within the constraint that they can only combine existing representations through weighted averages. The generation of precise computational results must ultimately occur through the feed-forward networks, which apply position-wise transformations to produce new content.
Architectural Impossibility of Exact Computation.
Feed-forward networks in transformers implement an expand-compress architecture first proposed in Vaswani et al. (2017). For arithmetic queries, the required computational results aren’t directly available in the attention-processed representations and must be computed—this is where the FFN serves as the primary computational engine.
This architecture can be understood as an implementation of the Universal Approximation Theorem (UAT) Hornik (1991): expanding to a higher-dimensional space (e.g., 49,152 dimensions in GPT-3) provides the width needed for universal function approximation, before compressing back to the model’s working dimensions (12,288). The fundamental operation follows:
$\text{FFN}(x)=\text{ReLU}(xW_{1}+b_{1})W_{2}+b_{2}$ (3)
Despite this theoretically universal approximation capacity, FFNs face a fundamental constraint for symbolic operations. The ReLU-based architecture implements piecewise linear functions, and as we prove formally in Appendix A for the case of multiplication, piecewise linear functions cannot implement exact symbolic operations requiring higher-order interactions through weight configuration alone. Similar constraints apply to division, exponentiation, and logical operations that require non-linear variable interactions over unbounded domains—while FFNs can approximate such operations arbitrarily well within any bounded region, symbolic computation requires exact results across potentially unbounded inputs, creating a fundamental architectural constraint that the expand-compress design cannot overcome through scale alone.
This impossibility extends to other exact symbolic operations, creating a fundamental architectural constraint that the expand-compress design cannot overcome through scale alone.
@@
Each layer follows the same computational pattern where attention processes the evolving representations and FFN generates the residual contribution. For layer $l$, the computation proceeds as:
$h_{\text{attn}}^{(l)}$ $=x^{(l-1)}+\text{Attention}_{l}(x^{(l-1)})$ (4)
$\Delta h^{(l)}$ $=\text{FFN}_{l}(h_{\text{attn}}^{(l)})$ (5)
$x^{(l)}$ $=x^{(l-1)}+\Delta h^{(l)}$ (6)
The key insight is that each layer’s FFN faces the identical architectural constraint—none can perform exact symbolic computation. Instead, they learn to coordinate their residual contributions toward the target output through hierarchical pattern decomposition:
  • Layer 1: Attention recognizes basic task structure (‘×’ indicates multiplication, operands ‘23’ and ‘17’), FFN contributes initial pattern fragment $\Delta h^{(1)}$ that begins moving toward ‘391’
  • Layer 2: Attention processes representations enriched by Layer 1’s computational work, FFN refines the approximation with $\Delta h^{(2)}$ based on learned pattern fragments
  • Layer L: Attention integrates all accumulated computational history, FFN makes final pattern-based adjustment $\Delta h^{(L)}$ toward the target
The final output becomes:
$\text{output}=\text{embeddings}+\sum_{l=1}^{L}\Delta h^{(l)}$ (7)
Critically, each FFN learns conditional pattern matching over attention-processed representations: “when attention gives me this enriched computational state $\to$ I output this residual contribution.” No individual layer performs algorithmic computation—each contributes learned pattern fragments that collectively approximate the target through residual accumulation.
Since exact algorithmic implementation remains architecturally impossible at each layer, gradient descent naturally converges toward hierarchical pattern assembly across the residual stream. This explains why models achieve high accuracy on familiar arithmetic patterns while failing systematically when problem structures deviate from training distributions—the system approximates multiplication through coordinated pattern retrieval rather than principled computation, leading to brittle performance on novel configurations with drops of up to 80% Srivastava et al. (2024); Mirzadeh et al. (2024).
@@
To test this hierarchical pattern decomposition hypothesis, we analyzed layer-by-layer hidden state trajectories in LLaMA-2-7B during arithmetic computation [^1] Code and data for reproducing all embedding analyses and layer-by-layer convergence experiments will be made available upon publication.. For prompts like “23 × 96 =” and “63.7 + 3.5 =”, we measured how each layer’s representation at the prediction position converges toward the final layer’s output representation.
@@
This pattern storage behavior aligns with recent findings by Dziri et al. Dziri et al. (2023), who demonstrate that transformers memorize computation patterns from training rather than learning generalizable algorithms, with models succeeding primarily when test patterns appeared frequently in training data. This frequency-dependent behavior also explains some relational reasoning failures we examine in Section 4.
But this raises a puzzling question: if LLMs cannot execute symbolic operations reliably, why do they excel at explaining algorithmic procedures? When asked “How do you multiply two numbers?", models provide textbook-perfect explanations of the algorithm. This suggests a potential solution: perhaps these articulated procedures could implicitly guide execution, with the model’s comprehensive knowledge of algorithms compensating for its execution limitations.
Unfortunately, this intuitive hope fails due to a third architectural constraint that prevents instructional knowledge from rescuing symbolic computation.
@@ 3.3 Claim 3: Next-Token Prediction Decouples Instruction from Execution
At inference time, different prompts steer the model toward geometrically distinct subspaces in its representational space. When prompted for algorithmic explanation (e.g., “How do you multiply two numbers?"), the model gravitates toward the instructional subspace, retrieving pedagogical patterns from its training distribution. When prompted for calculation (e.g., “What is 56 × 76?"), the model seeks out the execution subspace, accessing stored arithmetic patterns.
@@ 3.4 The Broader Pattern: From Architecture to Cognition
Recent detailed mechanistic analysis confirms this training-dependency hypothesis. Nikankin et al. Nikankin et al. (2024) provide valuable evidence that arithmetic computation relies on distributed “bags of heuristics” spread across multiple layers, rather than implementing dedicated algorithmic circuits for basic operations ($+$, --, $\times$, $\div$). Their neuron-level analysis reveals that no coherent computational circuits exist for these fundamental operations. Importantly, they find that models within the same training lineage (Llama3-8B vs Llama3-70B) share more similar heuristic patterns than models from different training backgrounds (Pythia-6.9B, GPT-J), yet even within this shared lineage, the models differ in the degree and sophistication of these patterns. This training-dependency gradient—where shared training yields similar but not identical mechanisms—confirms that apparent “arithmetic mechanisms” reflect training-specific statistical regularities rather than universal computational principles.

### v2

@@ 2.4 Compensatory Strategies and Architectural Remedies
The effectiveness of compensatory strategies has theoretical grounding in computational complexity. Li et al. (2024b) prove that chain-of-thought fundamentally changes the computational expressiveness of transformers: while constant-depth transformers without CoT can only solve problems in $\mathsf{TC}^{0}$, adding $T$ steps of CoT enables them to solve any problem solvable by Boolean circuits of size $T$. This provides the theoretical foundation for why self-scaffolding and chain-of-thought prompting partially succeed—they effectively transform parallel computation constraints into serial computation opportunities, allowing transformers to simulate $\mathsf{P/poly}$ with polynomially many intermediate steps. This theoretical result directly explains the empirical success of self-scaffolding as one of the compensatory strategies.
However, Peng et al. (2024) demonstrate that chain-of-thought requires an exponentially growing number of intermediate steps for complex function composition tasks, suggesting that self-scaffolding becomes ineffective when problems demand impractically long reasoning chains. This theoretical prediction aligns with empirical failures we discuss in Section 5.
@@
Tool-augmented approaches explicitly address the execution gap. Toolformer (Schick et al., 2023) enables automated tool calling, while ReAct (Yao et al., 2022) combines reasoning with external action. Yang et al. (2024) provide empirical validation that this separation—LLMs for predicate extraction, external systems for computation—consistently outperforms end-to-end LLM execution across arithmetic tasks.
@@ 3 Structural Limits on Symbolic Computation
The computational split-brain syndrome manifests most clearly in symbolic computation, where LLMs can perfectly explain algorithms they cannot reliably execute. This systematic dissociation—drawing analogy to neurological conditions where different brain systems cannot coordinate effectively—emerges as geometric separation in representational space: models develop distinct pathways for “knowing about” procedures versus “executing” them.
@@ 3.1 Claim 1: Contextual Averaging Prevents Automatic Domain Binding
Equation 1 is the familiar next-token prediction loss, where $\theta$ denotes the model parameters, $w_{t}$ is the token at position $t$, $w_{<t}$ represents all tokens before position $t$, and $p_{\theta}(\cdot|\cdot)$ is the conditional probability distribution learned by the model. Equation 2 shows how individual token embeddings are optimized: $e^{*}(w)$ is the optimal embedding for token $w$, $C(w)$ denotes all contexts containing token $w$, $e$ is the candidate embedding being optimized, $E(c\setminus w)$ represents embeddings of all tokens in context $c$ except $w$, and $w^{\prime}$ is the next token sampled from the data distribution $p_{\text{data}}(\cdot|c,w)$. This optimization produces context-averaged embeddings that trade off among multiple usage modes, similar to distributional semantics models such as Word2Vec (Mikolov et al., 2013).
@@
These geometric failures directly impede the formation of stable arithmetic circuits that could operate reliably across different input formats or numerical ranges. For LLMs to form stable, input-agnostic circuits for any algorithm, inputs must preserve the properties that algorithm expects, even when cast in high dimensions. Consider sorting—it fails if vector representations have random perturbations that violate ordering relationships. Without isometric properties in embedding space (preserving numerical order and relationships), models cannot form circuits that work consistently across all inputs. McLeish et al. (2024) demonstrate that altering input geometry alone can dramatically improve simple numerical reasoning, which might appear to contradict our contextual averaging argument. However, their approach involves training models exclusively on arithmetic problems, which inadvertently avoids the fundamental contextual averaging problem faced by real-world LLMs. Once embeddings are contaminated through multi-domain training—where tokens like “9.11” accumulate conflicting associations across historical, versioning, and mathematical contexts—this geometric solution breaks down. The McLeish results thus support rather than refute our analysis: they show embedding geometry matters precisely because real-world training makes it unsolvable.
@@ 3.2 Claim 2: Feed-Forward Networks Resort to Pattern Storage
Even if we solved the representation problem—imagine LLMs had perfect, mathematically-structured embeddings—would transformers be able to execute symbolic computation? The answer reveals a profound gap between capacity and capability. Feed-forward networks possess, through the Universal Approximation Theorem, the capacity to approximate any continuous function. Yet in practice, they systematically fail at symbolic computation, resorting instead to pattern storage. This paradox lies at the heart of the computational split-brain syndrome.
The Architectural Division of Labor.
To understand this paradox, we must first examine how computation flows through the transformer architecture and identify where symbolic computation must occur.
$\Delta h^{(l)}_{\text{attn}}$ $=\text{Attention}_{l}(x^{(l-1)})$ (3)
$h^{(l)}_{\text{attn}}$ $=x^{(l-1)}+\Delta h^{(l)}_{\text{attn}}$ (4)
$\Delta h^{(l)}_{\text{FFN}}$ $=\text{FFN}_{l}(h^{(l)}_{\text{attn}})$ (5)
$x^{(l)}$ $=h^{(l)}_{\text{attn}}+\Delta h^{(l)}_{\text{FFN}}$ (6)
$=x^{(l-1)}+\Delta h^{(l)}_{\text{attn}}+\Delta h^{(l)}_{\text{FFN}}$ (7)
Final output after $L$ layers:
$x^{(L)}=\text{embeddings}+\sum_{l=1}^{L}\Delta h^{(l)}_{\text{attn}}+\sum_{l=1}^{L}\Delta h^{(l)}_{\text{FFN}}$
Figure 2: Computational flow through transformer layers. When processing arithmetic operations (e.g., “43 × 78 = ?”), attention aggregates the operation context (operator and operands) while FFNs must generate the result. Due to causal masking, attention at the “=” position can only compute weighted averages of the embeddings it sees, creating a representation that cannot contain the novel direction for the result token.
Figure 2 illustrates the computational pipeline of transformer layers. [^2] We omit layer normalization and dropout from our analysis as these operations—being element-wise rescaling and masking—do not alter the fundamental computational limitations we identify. LayerNorm normalizes representations but cannot enable symbolic computation. Each transformer layer implements two successive residual connections: attention reads from the residual stream and adds its output back, creating an intermediate state that FFN then reads, transforms, and adds to produce the next layer’s representation.
Consider a concrete example: ‘43 × 78 = ?’. Due to causal masking at the ‘=’ position, attention only sees $\{43,\times,78,=\}$ and computes weighted averages of these embeddings. This creates a rich representation encoding ‘multiply 43 by 78’—identifying the operation and operands—but crucially, since attention can only compute weighted combinations of its input embeddings, this representation cannot contain the novel embedding direction needed for ‘3354’, which lies outside the space spanned by the input tokens.
Feed-forward networks then receive this attention-enriched representation and must somehow produce the correct result. Since attention cannot generate novel values, FFNs bear the entire computational burden: transforming the encoding of ‘multiply 43 by 78’ into the embedding direction for ‘3354’.
Theoretical Capacity via Universal Approximation.
In principle, FFNs have sufficient capacity for this task. Each FFN implements an expand-compress architecture (Vaswani et al., 2017):
$\text{FFN}(x)=W_{2}\cdot\text{ReLU}(W_{1}x+b_{1})+b_{2}$ (8)
By the Universal Approximation Theorem (Hornik, 1991), such networks can approximate any continuous function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ to arbitrary precision on compact sets. Theoretically, this includes multiplication: given sufficient width and appropriate weights, an FFN could approximate $x\times y$ within any bounded domain.
This raises a critical question: If FFNs have the capacity to approximate multiplication, why do they fail so catastrophically in practice?
The Training Reality: Role Specialization.
The answer lies not in what FFNs can do, but in what gradient descent trains them to do. We now present our central theoretical result:
Theorem (Role Specialization under Next-Token Prediction) Under gradient descent optimization for next-token prediction loss $\mathcal{L}=-\log p_{\theta}(w_{t}|w_{<t})$, the three parameter sets necessarily specialize into distinct computational roles: $\theta_{E}$ (embeddings) $\rightarrow$ Context-averaged representations $\theta_{A}$ (attention) $\rightarrow$ Operation context encoding $\theta_{F}$ (FFNs) $\rightarrow$ Memorized pattern-to-result mapping Proof Sketch: The cross-entropy loss requires $x^{(L)}\cdot e_{\text{result}}$ to be maximal among all vocabulary tokens. Since attention is constrained to weighted averages of inputs, it cannot generate an embedding direction for the result token that lies outside the space spanned by the input embeddings. FFNs must therefore learn the discrete mapping from attention’s operation encoding to result embeddings. While UAT guarantees this mapping can be approximated on the training set, gradient descent optimizes for memorizing specific pattern-to-token mappings rather than discovering algorithmic rules. See Appendix B for detailed analysis.
@@
To understand why memorization emerges inevitably, consider the optimization dynamics for arithmetic like ‘43 × 78 = 3354’. The gradient descent must minimize:
$\mathcal{L}=-\log p(\text{`3354'}|\text{`43 \texttimes 78 ='})$ (9)
This requires the final representation $x^{(L)}$ to have maximum dot product with the embedding of token ‘3354’ compared to all other vocabulary tokens. The gradients flow through three components with distinct constraints:
  1. 1.
    Embeddings lack the isometric properties required for arithmetic—numerical tokens cannot preserve ordering or distance relationships due to contextual averaging (Section 3.1).
  2. 2.
    Attention can only compute weighted averages of existing embeddings, unable to generate novel token directions outside their span.
  3. 3.
    FFNs must bridge the gap: mapping attention’s encoding to the specific embedding direction that maximizes probability for ‘3354’.
Since FFNs cannot implement true multiplication (piecewise linear functions cannot compute $x\times y$ exactly—see Appendix A), gradient descent converges to the only viable solution: memorizing a lookup table of pattern $\rightarrow$ result mappings. The UAT capacity that could theoretically approximate multiplication is instead allocated to storing these discrete mappings.
@@
To test our theoretical framework, we analyzed layer-by-layer hidden state trajectories in LLaMA-2-7B during arithmetic computation [^3] Code and data for reproducing all embedding analyses and layer-by-layer convergence experiments will be made available upon publication.. For prompts like “23 × 96 =” and “63.7 + 3.5 =”, we measured how each layer’s representation at the prediction position converges toward the final layer’s output representation.
@@
This pattern storage behavior aligns with recent findings by Dziri et al. (2023), who demonstrate that transformers memorize computation patterns from training rather than learning generalizable algorithms, with models succeeding primarily when test patterns appeared frequently in training data.
The Approximation-Computation Gap.
The fundamental distinction between what FFNs can do in theory versus practice reveals why pattern completion differs from algorithmic approximation:
  • UAT Approximation: FFNs can learn $f:K\to\mathbb{R}^{d}$ on compact training set $K$, successfully memorizing patterns like ‘23×17 $\rightarrow$ 391’.
  • Symbolic Computation: True algorithms execute for any valid input, maintaining systematic behavior on novel combinations.
  • The Key Difference: Robustness to distribution shift—algorithmic implementation (even imperfect) generalizes systematically, while pattern completion catastrophically fails outside memorized examples.
For novel inputs outside the training support $K$, FFNs have no algorithm to execute—only nearby memorized patterns to interpolate between. The discrete nature of token prediction exacerbates this: small errors in $x^{(L)}$ cause complete failure (wrong token), unlike continuous approximation where errors degrade gracefully. If students memorized multiplication tables then tried 3-digit problems, they would fail completely—only true algorithmic understanding enables generalization.
Many researchers, including ourselves initially, searched for generalizable arithmetic circuitry in LLMs. At inference time, such circuitry does not exist, as confirmed by studies showing LLMs rely on “bags of heuristics” rather than systematic algorithms (Nikankin et al., 2024; Tigges et al., 2024). Models perform pattern completion whether explaining algorithms or attempting execution—they just complete different memorized patterns.
Complexity-Theoretic Confirmation.
This empirical limitation aligns with theoretical impossibility results. Li et al. (2024b) prove that constant-depth transformers can only compute functions in $\mathsf{TC}^{0}$, while multiplication requires serial carry propagation—fundamentally outside $\mathsf{TC}^{0}$. Even with universal approximation capacity, architectural constraints prevent true symbolic computation in a single forward pass. The $\mathsf{TC}^{0}$ result establishes that no clever weight configuration can enable exact multiplication, while our analysis shows that training dynamics allocate the available approximation capacity toward pattern memorization rather than algorithmic discovery.
This analysis of pattern storage versus algorithmic computation raises a puzzling question: if LLMs cannot execute symbolic operations reliably, why do they excel at explaining algorithmic procedures? When asked “How do you multiply two numbers?”, models provide textbook-perfect explanations of the algorithm. This suggests a potential solution: perhaps these articulated procedures could implicitly guide execution, with the model’s comprehensive knowledge of algorithms compensating for its execution limitations. Unfortunately, this intuitive hope fails due to a third architectural constraint that prevents instructional knowledge from rescuing symbolic computation.
@@ 3.3 Claim 3: Next-Token Prediction Decouples Instruction from Execution
At inference time, different prompts steer the model toward geometrically distinct subspaces in its representational space. When prompted for algorithmic explanation (e.g., “How do you multiply two numbers?”), the model gravitates toward the instructional subspace, retrieving pedagogical patterns from its training distribution. When prompted for calculation (e.g., “What is 56 × 76?”), the model seeks out the execution subspace, accessing stored arithmetic patterns.
@@
3.4 Real-World Manifestation: Structured Data Operations
The architectural constraints we identify manifest concretely in real-world applications where organizations attempt to use LLMs for data analysis tasks. Consider a seemingly simple query: finding the maximum sales value across branches from a CSV table embedded in a prompt. This operation, trivial for any database system, requires:
  • Sequential iteration through records
  • Maintaining comparison state across rows
  • Tracking both the maximum value and its associated entity
  • Consistent numerical comparison operations
Recent empirical work by Wolff and Hulsebos (2025) provides direct evidence of these failures. Their evaluation of LLMs on analytical aggregations—including operations as fundamental as computing averages, sums, and maximum values—reveals “significant deficits in tabular reasoning performance.” These multi-step algorithms exemplify precisely the symbolic operations our three architectural constraints predict to be impossible through pattern completion alone.
The failure is systematic: MAX requires maintaining comparison state across iterations, AVG demands accumulating sums while counting elements, and even COUNT necessitates consistent increment operations. Each violates our identified constraints: contextual averaging (Claim 1) prevents stable numerical comparison, FFNs cannot implement exact iterative operations (Claim 2), and the instruction-execution disconnect (Claim 3) means that even when models can describe these algorithms perfectly, they cannot execute them reliably.
This has immediate practical implications. This has immediate practical implications. The inevitable failures are not due to inadequate training on tabular data or poor prompt engineering—they reflect fundamental architectural impossibilities that practitioners do not expect. As Wolff and Hulsebos (2025) demonstrate, performance degrades further with realistic data characteristics like missing values and duplicates, confirming these are architectural rather than training limitations.
@@ 3.5 The Broader Pattern: From Architecture to Cognition
Systematic vs. Pattern-Based Approximation.
Even when traditional systems use approximation, as in floating-point arithmetic, they maintain systematic behavior with guaranteed error bounds (IEEE 754). Each operation follows precise rules with predictable maximum errors. LLMs, by contrast, exhibit unsystematic pattern-based failures: correctly computing most arithmetic while failing unpredictably on specific cases, with no principled error bounds. For queries involving arithmetic operations, LLMs will therefore always encounter out-of-distribution problems if left to their own devices. This distinction highlights the fundamental difference between engineered approximation algorithms and emergent pattern retrieval.
@@
Recent detailed mechanistic analysis confirms this training-dependency hypothesis. Nikankin et al. (2024) provide valuable evidence that arithmetic computation relies on distributed “bags of heuristics” spread across multiple layers, rather than implementing dedicated algorithmic circuits for basic operations (++, -, $\times$, $\div$). Their neuron-level analysis reveals that no coherent computational circuits exist for these fundamental operations. Importantly, they find that models within the same training lineage (Llama3-8B vs Llama3-70B) share more similar heuristic patterns than models from different training backgrounds (Pythia-6.9B, GPT-J), yet even within this shared lineage, the models differ in the degree and sophistication of these patterns. This training-dependency gradient—where shared training yields similar but not identical mechanisms—confirms that apparent “arithmetic mechanisms” reflect training-specific statistical regularities rather than universal computational principles.
@@
Testable Predictions.
Our framework makes specific predictions that can be empirically validated:
  1. 1.
    Models trained exclusively on arithmetic should show different embedding geometry than general-purpose models, with more isometric numerical relationships
  2. 2.
    Pattern assembly signatures (gradual layer-by-layer convergence) should appear across all transformer model families
  3. 3.
    Instruction-execution separation should persist across architectures, including instruction-tuned and reasoning models
  4. 4.
    Novel format generalization should consistently fail while familiar format variations succeed (the key discriminator between pattern completion and algorithmic approximation)
These predictions provide a roadmap for future quantitative validation of our claims.